OpenStudy (anonymous):

f(x) = 3 + x^2 + tan (pix/2)

OpenStudy (anonymous):

Any one can help me with this derivative of an inverse

OpenStudy (sweetburger):

wait so your looking to take the derivative of this equation or what exactly?

OpenStudy (anonymous):

yes Im in the last part but I dont get quite the exact answer

OpenStudy (anonymous):

I get (pi sec^2 pix/2)/(2) + 2x

OpenStudy (freckles):

why do you say derivative of inverse ?

OpenStudy (freckles):

Do you mean just find derivative?

OpenStudy (sweetburger):

so 3 gets deleted x^2 turns into 2x and tan(pix/2) = sec^2x(pix/2) times the derivative of pix/2 which i believe equates to pi/2. I am not sure if this is totally correct so ya check with someone else.

OpenStudy (sweetburger):

so it would look like 2x+sec^2x(pix/2)(pi/2) I think

OpenStudy (anonymous):

yes, i evaluated at 0 but i get pi^2/4

OpenStudy (sweetburger):

x+sec^2(pix/2)(pi/2) is the correct value I accidentally left in x above.

OpenStudy (freckles):

I'm still confused why you said inverse

OpenStudy (freckles):

are you wanting to find \[(f^{-1})'(a)\] for some number a

OpenStudy (anonymous):

yes freckles

OpenStudy (freckles):

what is the a

OpenStudy (anonymous):

I used a couple of engines and I get (pi*sec^2(pix/2)(2))/ + 2x

OpenStudy (anonymous):

a = 3

OpenStudy (sweetburger):

@Eco1 you you are right I made a typo and left out the 2 in front of the x

OpenStudy (anonymous):

So I evaluate what I wrote above at zero but I dont get pi/2, which is the right asnwer

OpenStudy (freckles):

\[(f^{-1})'(3)=\frac{1}{f'(f^{-1}(3))} \\ f^{-1}(3)=? \implies f(?)=3 \\ \text{ so we need \to find ? such that }3=3+?^2+\tan(\frac{\pi ?}{2}) \\ \text{ and yes } f'(x)=0+2x+\frac{\pi}{2} \sec^2(\frac{\pi}{2}x)\]

OpenStudy (freckles):

that should be pretty easy to find ?

OpenStudy (freckles):

because we know both tan(0)=0 and 0^2=0

OpenStudy (freckles):

so ?=0

OpenStudy (freckles):

\[(f^{-1})'(3)=\frac{1}{f'(0)}\]

OpenStudy (freckles):

last thing for you to do is just to plug in 0 into your f' you found

OpenStudy (anonymous):

I did multiple times but I dont get the right answer

OpenStudy (freckles):

\[f'(0)=2(0)+\frac{\pi}{2}[\sec(\frac{\pi}{2}(0))]^2\]

OpenStudy (freckles):

well first the inside of that sec( ) thing you can be simplified pi/2*0=0 so sec(0)=?

OpenStudy (freckles):

\[(f^{-1})'(3)=\frac{1}{f'(0)}=\frac{1}{2(0)+\frac{\pi}{2}[\sec(0)]^2}\]

OpenStudy (freckles):

you know cos(0)=1 so sec(0)=?

OpenStudy (freckles):

hint sec and cos are reciprocal functions of each other

OpenStudy (anonymous):

reciprocal is synonym of inverse?

OpenStudy (freckles):

it depends what inverse you are talking about multiplicative inverse yeah

OpenStudy (freckles):

like sec(x)=1/cos(x) so sec(0)=1/cos(0)

OpenStudy (freckles):

1/cos(x) is reciprocal of cos(x)

OpenStudy (freckles):

some people just like to call it the flipping of

OpenStudy (anonymous):

but it is sec(0)^2 not sec(0)

OpenStudy (anonymous):

I meant sec^2(0)

OpenStudy (freckles):

if cos(0)=1 then sec(0)=1/cos(0)=1/1=1 if you square both sides of cos(0)=1 you get still cos^2(0)=1 so sec^2(0)=1 \[(f^{-1})'(3)=\frac{1}{f'(0)}=\frac{1}{0+\frac{\pi}{2}[1]^2}=\frac{1}{\frac{\pi}{2}}\]

OpenStudy (freckles):

1/(a/b) means just flip the number in the denominator

OpenStudy (freckles):

so 1/(a/b) is b/a

OpenStudy (freckles):

are you cool?

OpenStudy (freckles):

still processing?

OpenStudy (anonymous):

Thanks man, or girl?

OpenStudy (anonymous):

I just made an algebra error

OpenStudy (anonymous):

I got it now, but this problem is a pain in the behind

OpenStudy (freckles):

I might be a she. And also it gets way more fun with more practice.

OpenStudy (anonymous):

I think my algebra is rusty

OpenStudy (freckles):

it seems like everyone is working out their algebra kinks in calculus

OpenStudy (freckles):

so no worries

OpenStudy (anonymous):

tell me about it

OpenStudy (anonymous):

Any one out there who can give me a hand doing integrals

OpenStudy (freckles):

post a new question so people can see it :)